# Counterfactual Simulatability of LLM Explanations for Generation Tasks

### v1

@@ Abstract
LLMs can be unpredictable, as even slight alterations to the prompt can cause the output to change in unexpected ways. Thus, the ability of models to accurately explain their behavior is critical, especially in high-stakes settings. One approach for evaluating explanations is counterfactual simulatability, how well an explanation allows users to infer the model’s output on related counterfactuals. Counterfactual simulatability has been previously studied for yes/no question answering tasks. We provide a general framework for extending this method to generation tasks, using news summarization and medical suggestion as example use cases. We find that while LLM explanations do enable users to better predict LLM outputs on counterfactuals in the summarization setting, there is significant room for improvement for medical suggestion. Furthermore, our results suggest that the evaluation for counterfactual simulatability may be more appropriate for skill-based tasks as opposed to knowledge-based tasks.
@@ 1 Introduction
To anticipate these risks, recent work has turned attention to evaluating the reliability of LLM explanations (Madsen et al., 2024; Turpin et al., 2023; Kunz & Kuhlmann, 2024). In particular, Chen et al. (2024) evaluates the counterfactual simulatability of LLM explanations in yes/no question answering settings. Counterfactual simulatability is a measure of how well a model’s explanation allows humans to correctly infer the model’s predictions on simulatable counterfactuals (unseen inputs where the explanation enables the user to confidently guess the model’s output). Furthermore, according to Chen et al. (2024), the counterfactual simulatability of an explanation can be decomposed into simulation generality, a measure of the diversity of simulatable counterfactuals and simulation precision, a measure of the proportion of these counterfactuals for which humans correctly infer the model’s output. Ideal model explanations should balance both generality and precision.
Figure 1: Our evaluation pipeline. Given a model’s explanation, an LLM is prompted to generate relevant counterfactuals (right) and decompose the explanation into atomic units (left). For each unit, a human annotator verifies whether the element appears in the counterfactual (simulatability) and the counterfactual output (precision).
@@
Counterfactual simulatability, however, is equally critical to generation tasks, where the larger space of possible outputs makes understanding a model’s decision process more challenging. To fill this gap, we formalize a framework for evaluating counterfactual simulatability in language generation tasks (figure 1). We first use LLMs to produce counterfactuals and decompose explanations into atomic units. Then, human annotators determine the simulatability and precision of each unit of a models’ explanations. Rather than evaluating the factual correctness of explanations, our framework measures the ability of LLMs to accurately describe their behavior in a generative setting. The framework evaluates whether LLMs’ explanations lead to reliable human mental models consistent with their outputs.
@@
We apply the framework to two tasks: news summarization using CNN/DM (Nallapati et al., 2016) and medical suggestion generation using the Taiwan e-Hospital Dataset (Chen et al., 2022). These specific tasks involve different trade-offs between generality and precision. Explanations for news summarization can be highly general as models may employ similar approaches towards news documents that contain common elements (e.g., dates, quotes) while differing in content. In contrast, medical suggestion explanations can be less general as the model’s suggestion may be highly dependent on the details described by the user (i.e., slight differences in symptoms can lead to completely different suggestions). This leads to a more limited set of counterfactuals the explanations apply to, though potentially being more precise in its outputs.
With these generation tasks, we conduct an initial evaluation of LLMs’ explanations and assess where models’ explanations in these generation tasks fail.
Our contributions are summarized as follow:
@@
    We assess the feasibility of using LLMs to automate human annotation in our evaluation pipeline and find that an LLM achieves significant agreement with humans.
@@
    Using our framework, we evaluate Chain-of-thought and Post-hoc explanations for multiple LLMs on two complementary tasks: news summarization and medical suggestion generation. We show that LLMs generate explanations that lead to reliable mental models in news summarization, but not medical suggestion.
@@ 2 Related work
Human mental models. Humans form mental models of the physical world as a whole (Gentner & Stevens, 2014) as well as specific technologies (e.g., Payne (1991); Du et al. (2018); Lei et al. (2016)) through their past experiences and observations. Specifically with regard to artificial intelligence systems, explanations of model predictions have been considered high quality if they provide users with an accurate and generalizable understanding of the system in the form of mental models (Rutjes et al., 2019; Merry et al., 2021). When such explanations are successful, they can improve users’ ability to effectively use AI models (Vasconcelos et al., 2022; Senoner et al., 2024) as well as to anticipate and correct undesirable model behaviors, such as biases and incorrect predictions (Bansal et al., 2019). We specifically evaluate LLMs’ explanations in generation tasks considering their ability to help form mental models.
Explanation evaluation. A variety of different approaches have been used to evaluate the quality and utility of model-generated natural language explanations and rationales. Studies evaluating natural language explanations similar to word attributions (Huang et al., 2023; Madsen et al., 2024) as well as unconstrained explanations (Turpin et al., 2023) have focused on faithfulness measures. In particular, work has proposed metrics for dimensions including comprehensiveness (DeYoung et al., 2020), sufficiency (DeYoung et al., 2020), alignment with human rationales (Fayyaz et al., 2024), and scrutability (Xu et al., 2023) among others. In contrast to intrinsic measures of model explanations, other work evaluates explanations extrinsically through impact on task performance (e.g., Camburu et al. (2018); Wei et al. (2022); Krishna et al. (2023)). Among these topics, relatively few works have investigated natural language explanations in language generation tasks; such studies include dialogue responses (Zhou et al., 2021), dialogue understanding (Gao et al., 2024), and more prominently, open-ended question answering (Ho et al., 2023; Fragkathoulas & Chlapanis, 2024; Lyu et al., 2023). While Chen et al. (2024) introduces counterfactual simulatability in classification settings, we propose a novel framework for counterfactual simulatability in language generation settings.
@@ 3 Counterfactual simulatability for generation tasks
Explanations can be evaluated by considering to what extent an observer, having seen a model’s explanation for some input, can infer (i.e., simulate) the model’s output for a counterfactual input. For instance, if a user asks an LLM “do dolphins swim?” and a model answers “yes” with the explanation “all aquatic animals swim”, then the user would infer that when asked the counterfactual “do starfish swim?”, the model will similarly answer “yes”. If, in reality, the model answers “no” to this question, then, as Chen et al. (2024) notes, the explanation is ineffective because it creates a mental model that is inconsistent with the model’s behavior despite the answer being factually correct. Counterfactual simulatability measures how accurate these mental models are on simulatable counterfactuals, unseen inputs where the explanation should allow the user to confidently guess the model’s output (Chen et al., 2024).
However, in contrast to this classification example, generation tasks have a much larger output space (e.g., there exists many possible summaries for a news document or free text responses to an open-ended medical question). This makes simulation extremely challenging as it is nearly impossible to precisely identify a single possible output based on its explanation. For instance in news summarization, if a user is shown the explanation “the summary should include the key event” and a counterfactual document on “the opening ceremony of the 2024 Paris Olympics”, while the user can logically infer that the summary will include the opening ceremony, it is impossible to predict the exact wording in which it will appear. Explanations typically cannot enable humans to pinpoint a single model output. However, they are still very useful if they help humans narrow down the possible outputs (e.g., refining “all possible summaries” to “summaries that mention the opening ceremony”).
@@ 3.2 Simulatability
Given a mental model, a counterfactual input is deemed simulatable if the observer can refine their expectation of the model’s output on the counterfactual (i.e., $|h_{x,e_{x}}(x^{\prime})|\ll O$). Our approach to determining whether a counterfactual is simulatable is outlined further in section 4.3.
3.3 Simulation generality and precision
The generality of an explanation is defined as the diversity of simulatable counterfactuals it applies to. In the previous example, the explanation “the summary should include the key event” is more general compared to “the summary should mention the opening ceremony of the 2024 Paris Olympics” since the former can be applied to documents covering a broader range of topics. Generality is calculated as one minus the average pairwise similarity between simulatable counterfactuals:
$\text{generality}=1-\mathbb{E}_{x^{\prime},x^{\prime\prime}\sim p,x^{\prime}\neq x^{\prime\prime}}[\alpha(x^{\prime},x^{\prime\prime})],$ (1)
where $p$ is the distribution of simulatable counterfactuals and $\alpha$ is a similarity metric such as cosine similarity. Precision measures whether the model’s actual output $M(x^{\prime})$ is within the observer’s inferred output space:
$\text{precision}=\mathbb{E}_{x^{\prime}\sim p}\left[\mathbf{1}[M({x^{\prime}})\in h_{e_{x}}(x^{\prime})]\right].$ (2)
In practice, it is unfeasible for humans to enumerate all valid outputs in $h_{e_{x}}(x^{\prime})$ due to the large output space and so determining whether the model’s output is a member of the human’s inferred output space is difficult. In section 4.3, we propose a method to estimate mental models by evaluating the presence of inferred atomic units of information in the model’s output.
@@ 4.1 Datasets
We select two generative tasks that represent practical LLM use cases and whose natural language explanations differ in relation to generality and precision. For the first task of news summarization, we use the CNN/DM dataset (Nallapati et al., 2016). Summarization is chosen as it is a well-studied application for language generation. Additionally, this task allows us to study high-level, abstract explanations since summarization does not rely on facts learned during the LLM’s pre-training, rather its ability to identify key elements from the input document.
Conversely, we also investigate medical suggestion, a domain where faithful explanations are critical as they allow humans to accurately predict the model’s behavior. For this task, we use the Taiwan e-Hospital Dataset, a collection of 86,399 Mandarin question answer pairs from the online health website Taiwan e-Hospital (Chen et al., 2022). We select this specific dataset because each sample consists of a question, suggestion, explanation triple, ensuring that the questions are complex enough such that explanations are helpful. Furthermore, this dataset mirrors the practical setting we aimed to study, where everyday users interact with LLMs to seek general medical advice without any specialized knowledge.
The differences between explanations for these tasks are illustrated in figure 2. In contrast to summarization, generating medical suggestions is more knowledge-based, requiring the LLM to identify key elements of the user’s query (e.g. their expressed symptoms, any relevant medical history) and relate them to potential suggestions using knowledge encoded in the model. As such, the explanations are highly specific to the input. For instance in figure 2, whereas the summarization explanation applies to counterfactuals where abstract elements such as a “key event” or “quotes or states’ are mentioned, the medical suggestion explanation can only be used to infer the model’s output when a user expresses “persistent nasal discharge”, “frequent sneezing”, and “symptoms [that] have lasted for over three months”. While explanations for medical suggestion are likely to apply to a more limited range of counterfactuals, they allow the user to infer specific pieces of information in the output as opposed to high-level elements as is the case for summarization. Therefore, these two tasks are complementary for studying counterfactual simulatability as they represent different requirements for generality and precision.
@@ 4.2 Explanations
Through prompting (instructions, few-shot examples), we guide the LLMs to generate explanations that emphasize this difference in generality and precision across tasks. For news summarization, to be highly generalizable, we encourage the model to explain its decision process at a high-level, identifying abstract elements while avoiding reference to specific topics. For medical suggestion, we instruct the model to first identify important details in the user’s question, propose possible underlying causes, and suggest recommended actions accordingly. These approaches reflect the skill-based vs. knowledge-based nature of our tasks. Additionally, following Chen et al. (2024), we experiment with both Chain-of-thought and Post-hoc prompting to produce explanations (Camburu et al., 2018). The prompts used are provided in appendix F.
@@ 4.3 Estimating mental models
Generation tasks are particularly challenging for counterfactual simulatability because mental models are extremely difficult to estimate given the large space of possible outputs. To this end, we introduce a method for estimating mental models by considering atomic units of information in the explanation as a proxy.
Intuitively, explanations $e_{x}$ identify key pieces of information that the model considers essential for the output. Thus, an observer may simulate that if the counterfactual input $x^{\prime}$ includes a piece of information $a$ deemed important by the model in $e_{x}$, the model should similarly take into account $a$ in its output for $x^{\prime}$. Based on this explanation $e_{x}$, an observer’s mental model $h_{e_{x}}$ can be formalized as:
$h_{e_{x}}(x^{\prime})=\{o\in O\mid\forall a\in e_{x}\cap x^{\prime},a\in o\},$ (3)
where the space of possible outputs on the counterfactual is refined to only outputs that contain the elements $\{a\}$. With this, we formulate simulatability and precision as:
$\text{simulatability}=\mathbb{1}\{\forall a\in e_{x},a\in x^{\prime}\}$ (4)
@@
First, we determine whether a counterfactual is simulatable by verifying whether all atomic units of the explanation appear in the counterfactual (equation 4). Then, for simulatable counterfactuals, precision is measured as the proportion of these atomic units that appear in the model’s counterfactual output (equation 5). Generality is calculated using the approach outlined in section 3.3 using cosine similarity. Finally, we average generality and precision scores across explanations.
A key distinction between the tasks is that while each atomic unit of a summarization explanation is linked to an expected item in the input (for simulatability) and its reference in the output (for precision), medical suggestion explanations do not follow this one-to-one mapping of units between input and output. For instance, using the example provided in figure 2, while ”Persistent nasal discharge” is identified as important, it is unclear which suggested action the model has generated in response to this symptom. To address this, for the medical domain, we further instruct the LLM when parsing the explanation to classify units into two categories, patient information and suggestions. These categories are then used to evaluate simulatability and precision respectively during annotation (see figure 2).

### v2

@@ Abstract
LLMs can be unpredictable, as even slight alterations to the prompt can cause the output to change in unexpected ways. Thus, the ability of models to accurately explain their behavior is critical, especially in high-stakes settings. Counterfactual simulatability measures how well an explanation allows users to infer the model’s output on related counterfactuals and has been previously studied for yes/no question answering. We provide a general framework for extending this method to generation tasks, using news summarization and medical suggestion as example use cases. We find that while LLM explanations do enable users to better predict their outputs on counterfactuals in the summarization setting, there is significant room for improvement for medical suggestion. Furthermore, our results suggest that evaluating counterfactual simulatability may be more appropriate for skill-based tasks as opposed to knowledge-based tasks.
Counterfactual Simulatability of LLM Explanations for Generation Tasks
Marvin Limpijankit, Yanda Chen, Melanie Subbiah, Nicholas Deas, and Kathleen McKeown Department of Computer Science, Columbia University ml4431@columbia.edu
@@ 1 Introduction
To anticipate these risks, recent work has turned attention to evaluating the reliability of LLM explanations (Madsen et al., 2024; Turpin et al., 2023; Kunz and Kuhlmann, 2024). In particular, Chen et al. (2024) evaluates the counterfactual simulatability of LLM explanations for yes/no question answering. Counterfactual simulatability is a measure of how well a model’s explanation allows humans to correctly infer the model’s predictions on simulatable counterfactuals (unseen inputs where the explanation should enable the user to confidently guess the model’s output). Furthermore, according to Chen et al. (2024), the counterfactual simulatability of an explanation can be decomposed into simulation generality, a measure of the diversity of simulatable counterfactuals and simulation precision, a measure of the proportion of these counterfactuals for which humans correctly infer the model’s output. Ideal model explanations should balance both generality and precision.
@@
Counterfactual simulatability, however, is equally critical to generation tasks, where the larger space of possible outputs makes understanding a model’s decision process more challenging. To fill this gap, we formalize a framework for evaluating counterfactual simulatability in language generation tasks (Figure 1). To evaluate a model’s explanation, we first use a separate LLM to (1) decompose the explanation into atomic units and (2) generate relevant counterfactuals. Then, an annotator (human or the LLM) evaluates the simulatability and precision of each unit of the model’s explanation using the counterfactual and the model’s output on the counterfactual respectively. Finally, generality and precision scores are calculated following §4.4. Rather than evaluating the factual correctness of explanations, our framework measures the ability of LLMs to accurately describe their behavior in a generative setting. We evaluate whether LLMs’ explanations lead to reliable human mental models that are consistent with their outputs.
Figure 1: Our evaluation pipeline. Given a model’s explanation, an LLM is prompted to decompose the explanation into atomic units (left) and generate relevant counterfactuals (right). For each unit, an annotator verifies whether the element appears in the counterfactual (simulatability) and the counterfactual output (precision).
@@
We apply the framework to two tasks: news summarization using CNN/DM (Nallapati et al., 2016) and medical suggestion generation using the Taiwan e-Hospital Dataset (Chen et al., 2022b). These tasks involve different trade-offs between generality and precision. Explanations for news summarization can be highly general as models may employ similar approaches for documents that contain common elements (e.g., dates, quotes) but differ in content. In contrast, medical suggestion explanations may be less general as the model’s response can be highly dependent on the user input (i.e., minor variances in expressed symptoms can lead to different suggestions). This can lead to a more limited set of simulatable counterfactuals, though potentially being more precise in its outputs. With these generation tasks, we conduct an initial evaluation of LLMs’ explanations and assess where models’ explanations in these generation tasks fail.
In summary, our contributions are as follows:
@@
    We assess the feasibility of using an LLM to automate human annotation in our evaluation pipeline and find that doing so achieves agreement with humans comparable to agreement among human annotators.
@@
    Using our framework, we evaluate Chain-of-thought and Post-hoc explanations for multiple LLMs on two complementary tasks: news summarization and medical suggestion generation. We show that LLM explanations lead to reliable mental models in news summarization, but not medical suggestion.  [^1] Our code is available at github.com/mlimpijankit/counterfactual-simulatability-generation-tasks
@@ 2 Related Work
Human mental models. Humans form mental models of the physical world as a whole (Gentner and Stevens, 2014) as well as specific technologies (e.g., Payne, 1991; Du et al., 2018; Lei et al., 2016) through their past experiences and observations. Specifically with regard to artificial intelligence systems, explanations of model predictions have been considered high quality if they provide users with an accurate and generalizable understanding of the system in the form of mental models (Rutjes et al., 2019; Merry et al., 2021). When such explanations are successful, they can improve users’ ability to effectively use AI models (Vasconcelos et al., 2022; Senoner et al., 2024) as well as to anticipate and correct undesirable model behaviors, such as biases and incorrect predictions (Bansal et al., 2019). We specifically evaluate LLMs’ explanations in generation tasks considering their ability to help form mental models.
Explanation evaluation. A variety of different approaches have been used to evaluate the quality and utility of model-generated natural language explanations and rationales. Studies evaluating natural language explanations similar to word attributions (Huang et al., 2023; Madsen et al., 2024) as well as unconstrained explanations (Turpin et al., 2023) have focused on faithfulness measures. In particular, work has proposed metrics for dimensions including comprehensiveness (DeYoung et al., 2020), sufficiency (DeYoung et al., 2020), alignment with human rationales (Fayyaz et al., 2024), plausibility (Wojciechowski et al., 2024), and scrutability (Xu et al., 2023) among others. In contrast to intrinsic measures of model explanations, other work evaluates explanations extrinsically through impact on task performance (e.g., Camburu et al., 2018; Wei et al., 2022; Krishna et al., 2023). Among these topics, relatively few works have investigated natural language explanations in language generation tasks; such studies include dialogue responses (Zhou et al., 2021), dialogue understanding (Gao et al., 2024), and more prominently, open-ended question answering (Ho et al., 2023; Fragkathoulas and Chlapanis, 2024; Lyu et al., 2023). While Chen et al. (2024) introduces counterfactual simulatability in a binary classification setting, we propose a novel framework for counterfactual simulatability in language generation settings.
@@ 3 Counterfactual Simulatability for Generation Tasks
Explanations can be evaluated by considering whether an observer, having seen a model’s explanation for some input, can infer (i.e., simulate) the model’s output on a related counterfactual input. For instance, if a user asks “do dolphins swim?” and a model answers “yes” with the explanation “all aquatic animals swim”, then the user would infer that when asked the counterfactual “do starfish swim?”, the model will similarly answer “yes”. If, in reality, the model answers “no” to this question, then, as Chen et al. (2024) notes, the explanation is ineffective because it creates a mental model that is inconsistent with the model’s behavior, even though the answer may be factually correct. More specifically, counterfactual simulatability measures how accurate these mental models are on simulatable counterfactuals, unseen inputs where the explanation allows the user to confidently predict the model’s output (Chen et al., 2024).
However, in contrast to the classification example, generation tasks have a much larger output space (e.g., there exists many possible summaries for a news document). This makes simulation extremely challenging as it is impossible to precisely identify a single possible output based on an explanation. For instance in news summarization, if a user is shown the explanation “the summary should include the key event” and a counterfactual document on “the opening ceremony of the 2024 Paris Olympics”, while the user can logically infer that the summary will include the opening ceremony, it is impossible to predict the exact wording in which it will appear. Explanations typically cannot enable humans to pinpoint a single model output. However, they are still very useful if they help humans narrow down the possible outputs (e.g., refining “all possible summaries” to “summaries that mention the opening ceremony”).
@@ 3.2 Simulatability
Simulatability is a metric on counterfactuals that describes whether the model’s output on that counterfactual can be inferred given an explanation. More specifically, a simulatable counterfactual is one where an observer, using the mental model produced by an explanation, can refine their expectation of the model’s output (i.e., $|h_{x,e_{x}}(x^{\prime})|\ll O$). Non-simulatable counterfactuals are omitted from our evaluation as the explanations are not helpful in these cases.
3.3 Simulation generality
Simulation generality is a metric on explanations that measures the diversity of simulatable counterfactuals an explanation leads to. For instance, the explanation “the summary should include the key event” is more general compared to “the summary should mention the opening ceremony of the 2024 Paris Olympics” since the former can be applied to documents covering a broader range of topics. It is calculated as one minus the average pairwise similarity between simulatable counterfactuals:
$\text{generality}=1-\mathbb{E}_{x^{\prime},x^{\prime\prime}\sim p,x^{\prime}\neq x^{\prime\prime}}[\alpha(x^{\prime},x^{\prime\prime})]$ (1)
where $p$ is the distribution of simulatable counterfactuals and $\alpha$ is a similarity metric such as cosine similarity. Explanations with higher generality are more useful since they enable users to infer the model’s behavior on a wider range of possible inputs.
3.4 Simulation precision
Simulation precision is another metric on explanations and measures whether the model’s actual output $M(x^{\prime})$ is within the observer’s inferred output space:
$\text{precision}=\mathbb{E}_{x^{\prime}\sim p}\left[\mathbf{1}[M({x^{\prime}})\in h_{e_{x}}(x^{\prime})]\right]$ (2)
Explanations with higher precision scores indicate that the mental models they produce are more aligned with the model’s actual behavior. In practice, it is unfeasible for humans to enumerate all valid outputs in $h_{e_{x}}(x^{\prime})$ due to the large output space. Thus, in §4.4 we propose a method to estimate mental models by evaluating the atomic units of a model’s explanation.
@@ 4.1 Datasets
We select two generative tasks that represent practical LLM use cases and whose natural language explanations differ in generality and precision. For the first task of news summarization, we use the CNN/DM dataset (Nallapati et al., 2016). Summarization is chosen as it is a well-studied application for language generation. Additionally, this task allows us to study high-level, abstract explanations since summarization does not rely on facts learned during the LLM’s pre-training, rather its ability to extract key elements from the input document.
Conversely, we also investigate medical suggestion, a domain where explanations are critical. For this task, we use the Taiwan e-Hospital Dataset, a collection of 86,399 Mandarin question answer pairs from the online health website Taiwan e-Hospital (Chen et al., 2022b). The data was translated into English using the Google Translate API. We select this specific dataset because each sample consists of a question, suggestion, explanation triplet, ensuring that the questions are complex enough such that explanations are helpful. Furthermore, the dataset also mirrors the practical setting we aim to study, where everyday users interact with LLMs to seek general medical advice without any specialized knowledge.
The contrast between explanations for these tasks is illustrated in Figure 2. Medical suggestions are more knowledge-based, requiring the LLM to identify key aspects of the query (e.g. expressed symptoms, relevant medical history) and generate suggestions using knowledge encoded in the model. As such, the explanations are highly specific to the input content. In the example above, the summarization explanation applies to counterfactuals containing any “key event” or “quotes or statements’ whereas the medical suggestion explanation can only be used to infer the model’s output when a user expresses “persistent nasal discharge”, “frequent sneezing”, and “symptoms [that] have lasted for over three months”. While medical suggestion explanations likely apply to a more limited set of counterfactuals, they allow the user to infer specific pieces of information rather than high-level elements as is the case for summarization. Therefore, these tasks are complementary for studying counterfactual simulatability as they represent different requirements for generality and precision.
4.2 Models
Our proposed framework uses two models: the LLM being evaluated and the LLM used to automate parts of the pipeline (e.g., counterfactual generation, explanation decomposition, annotation). We evaluate Anthropic’s Claude 3.7 Sonnet, OpenAI’s GPT-4 and GPT-4 Turbo, and Meta’s Llama 3.3 70B Instruct. These represent a few popular proprietary and open-source models. To automate steps in the pipeline, we opt for GPT-4 Turbo due to its competitive performance at a low cost.
@@ 4.3 Explanations
Through prompting (instructions, few-shot examples), we guide the LLMs to generate explanations that emphasize the difference in generality and precision across tasks. For news summarization, to be highly generalizable, we encourage the model to explain its decision process at a high-level, identifying abstract elements while avoiding reference to specific topics. For medical suggestion, we instruct the model to first identify important details in the user’s question, propose possible underlying causes, and suggest recommended actions accordingly. These approaches reflect the skill-based vs. knowledge-based nature of our tasks. Additionally, following Chen et al. (2024), we experiment with both Chain-of-thought and Post-hoc prompting to produce explanations (Camburu et al., 2018). The prompts used are provided in Appendix D.
@@ 4.4 Estimating mental models
Generation tasks are challenging for counterfactual simulatability because mental models are extremely difficult to estimate given the large space of possible outputs. Motivated by previous work on atomic units (Wright et al., 2022; Chen et al., 2022a; Kamoi et al., 2023), we introduce a method for estimating mental models by decomposing explanations into units of information as a proxy.
Intuitively, explanations $e_{x}$ identify key pieces of information that the model considers essential for the output. Thus, an observer may simulate that if the counterfactual input $x^{\prime}$ includes a piece of information $a$ deemed important in $e_{x}$, the model should similarly take into account $a$ in its output for $x^{\prime}$. Based on this explanation $e_{x}$, an observer’s mental model $h_{e_{x}}$ can be formalized as:
$h_{e_{x}}(x^{\prime})=\{o\in O\mid\forall a\in e_{x}\cap x^{\prime},a\in o\}$ (3)
where the space of possible outputs on the counterfactual is refined to only outputs that contain the elements $\{a\}$. With this, we formulate simulatability, generality, and precision for each explanation, counterfactual pair as follows. First, the simulatability of a counterfactual is determined by verifying if all atomic units of the explanation appear in the counterfactual:
$\text{simulatability}=\mathds{1}\{\forall a\in e_{x},a\in x^{\prime}\}$ (4)
Then, all non-siumulatable counterfactuals are discarded and the generality of an explanation is calculated as one minus the average pairwise cosine similarity between simulatable counterfactuals (Equation 1). Finally, by identifying whether each unit of the explanation is addressed in the model’s output on the counterfactual, a precision score is calculated as:
@@
Note that this differs from the formulation in §3.4. Precision scores are averaged across counterfactuals for a given explanation.
Explanation Decomposition.
The atomic units of an explanation should highlight aspects that are considered important to the model, and thus, a user would expect in the output accordingly. These effectively describe a user’s mental model and should ideally be separate, non-overlapping points. One key distinction between the tasks is that while each atomic unit of a summarization explanation is linked to an expected item in the input (for simulatability) and its reference in the output (for precision), medical suggestion explanations do not follow this one-to-one mapping of units between input and output. For example in Figure 2, while ”Persistent nasal discharge” is identified as important, it is unclear which suggested action the model has generated in response to this symptom. To address this, for the medical domain, we further instruct the LLM when decomposing the explanation to classify units into two categories, patient information and suggestions. Atomic units from these categories are then used to evaluate simulatability and precision respectively during annotation (Figure 2).